# Variational Autoencoders for Efficient Simulation-Based Inference

### v1

@@ Abstract
We present a generative modeling approach based on the variational inference framework for likelihood-free simulation-based inference. The method leverages latent variables within variational autoencoders to efficiently estimate complex posterior distributions arising from stochastic simulations. We explore two variations of this approach distinguished by their treatment of the prior distribution. The first model adapts the prior based on observed data using a multivariate prior network, enhancing generalization across various posterior queries. In contrast, the second model utilizes a standard Gaussian prior, offering simplicity while still effectively capturing complex posterior distributions. We demonstrate the efficacy of these models on well-established benchmark problems, achieving results comparable to flow-based approaches while maintaining computational efficiency and scalability.
1 Introduction
@@
The simulation-based inference (SBI) problem involves statistical inference of parameters $\boldsymbol{\theta}$ of a stochastic simulation model $M(\boldsymbol{\theta},\xi)$ containing random states $\xi$, from observed data $\boldsymbol{y}$ (Cranmer et al., 2020). Once calibrated, the simulation model can be used to reason about, analyze and explain the observed data $\boldsymbol{y}$ in the context of the corresponding physical process (Lavin et al., 2021). Several challenges arise in SBI owing to model stochasticity and potential multi-valuedness, where different sets of parameter values can produce similar observations, or similar parameters may lead to varied outputs. Consequently, the task is to infer the posterior distribution $p(\boldsymbol{\theta}\mid\mathbf{y})$ in a Bayesian setting.
@@
A user-defined prior distribution $p(\boldsymbol{\theta})$ encodes domain knowledge or assumptions about plausible parameter values. The likelihood function $p(\mathbf{y}\mid\boldsymbol{\theta})$ is however, often intractable due to intricate stochastic dynamics, high-dimensional integration, or latent variables making analytical computation impractical (Sisson et al., 2018). Although a closed-form expression for the likelihood function $p(\mathbf{y}\mid\boldsymbol{\theta})$ is unavailable, samples from this distribution can be generated by evaluating the simulator $M$ with varying $\xi$ for different $\boldsymbol{\theta}$ drawn from the prior, effectively yielding samples from the joint distribution $(\boldsymbol{\theta},\mathbf{y})\sim p(\boldsymbol{\theta},\mathbf{y})$.
@@
Deep learning based approaches that directly approximate the model likelihood, or the posterior distribution have gained increasing popularity in the SBI setting (Zammit-Mangion et al., 2024). Several approaches involve flow-based models, which are potentially compute-intensive and also restrict the design space of neural networks to invertible functions with computationally efficient Jacobian calculations (Rezende and Mohamed, 2015). This constraint can limit the flexibility of the model in handling more complex data structures. Alternative approaches like Generative Adversarial Training for SBI (GATSBI) (Ramesh et al., 2022) use an adversarial network, where a generator produces simulations that mimic observed data, and a discriminator distinguishes between real and simulated data. However, the adversarial approach can suffer from training instability, mode collapse, and difficulties in balancing the generator and discriminator networks (Arjovsky and Bottou, 2022; Arjovsky et al., 2017), complicating training and limiting robustness.
@@
Contribution: We propose a variational inference approach for simulation-based inference, utilizing a Conditional Variational Autoencoder (C-VAE) architecture (Kingma and Welling, 2014; Sohn et al., 2015) that is simple, computational efficient, scalable, interpretable and robust. As VAEs are not constrained by the invertibility requirement, they allow seamless integration of various neural network architectures without significant modifications as opposed to flow-based models. VAEs also avoid the instability associated with GANs due to adversarial learning dynamics. As a result, C-VAEs can handle complex data structures and dependencies more effectively, making them suitable for a wider range of tasks without requiring external summary networks (Cranmer et al., 2020).
@@
The paper is organized as follows. Section 2 outlines the related work. In Section 3, we provide a formal problem definition, followed by a detailed explanation of the proposed architecture in Sections 3.1 and 3.2. A comprehensive comparison on benchmark test problems is presented in Section 4 and in the Appendix. Section 5 discusses the results, while Section 6 concludes the paper by summarizing the key findings and further advancements.
@@ 2 Related Work
A vast majority of SBI approaches based on deep learning aim to either approximate the likelihood (Hermans et al., 2020), or focus on directly approximating the posterior (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019). Additionally, there exist methods that target both the likelihood and posterior simultaneously (Radev et al., 2023a; Glöckler et al., 2022; Wiqvist et al., 2021)
@@
Neural Posterior Estimation (NPE) (Papamakarios and Murray, 2016) was one of the earliest approaches towards one-step posterior estimation, introducing the use of conditional normalizing flows to approximate the posterior distribution. By incorporating techniques such as Neural Spline Flows (NSFs) and Mixture Density Networks (MDNs), NPE enhanced both the accuracy and scalability of simulation-based inference. In comparison to traditional approaches like Approximate Bayesian Computation (ABC) or Sequential Monte-Carlo (SMC)-ABC (Sisson et al., 2018) that rely on repeated simulations and rejection-based schemes, NPE and its derivatives such as Robust NPE (Ward et al., 2022) are more efficient and flexible.
@@
Sequential Neural Posterior Estimation (SNPE) improves NPE’s efficiency by employing sequential training, where initial parameters are sampled from the prior, and the model is iteratively refined by focusing on regions of the parameter space that are most relevant to the parameter inference task. This adaptive sampling allows for more efficient exploration of the posterior distribution. Early versions like SNPE-A (Papamakarios and Murray, 2016) and SNPE-B (Lueckmann et al., 2017) required correction steps to account for changes in the sampling distribution, adding complexity to the process. In contrast, Automatic Posterior Transformation (APT) or SNPE-C (Greenberg et al., 2019), eliminated this need, offering a more robust and streamlined approach by leveraging normalizing flows to refine the posterior iteratively. However, a limitation of sequential methods like APT is that they necessitate multiple rounds of training, extending the total training time.
@@
The recently proposed Jointly Amortized Neural Approximation (JANA) framework (Radev et al., 2023a) addresses posterior and likelihood estimation through joint amortization. JANA employs conditional invertible neural networks (cINNs) for both the posterior and likelihood networks, allowing efficient transformations between the parameter space and latent variables. Additionally, JANA includes a trainable summary network sub-module optimized to extract maximally informative data representations in an end-to-end manner. Unlike previous approaches, JANA follows a fully amortized strategy, allowing the evaluation of normalized densities and generation of conditional random draws for parameter estimation and surrogate modeling. Simformer (Gloeckler et al., ) is a recent approach that leverages a probabilistic diffusion model with transformer architectures to accurately learn complex, high-dimensional posteriors, making it highly flexible. However, the transformer-based architecture is relatively resource-intensive and involves careful optimization.
@@
We compare our proposed approach with four popular SBI methods: JANA (Radev et al., 2023a), GATSBI (Ramesh et al., 2022), APT (Greenberg et al., 2019), and NPE (Papamakarios and Murray, 2016). The motivation is to cover a variety of approaches (e.g., flow-based, adversarial, one-shot, sequential) while noting that our focus is on efficiency and scalability as opposed to solely the posterior approximation accuracy. We refer the reader to (Zammit-Mangion et al., 2024) and (Radev et al., 2023a) for further reading on deep learning methods for SBI.
@@ 3 Simulation Based Inference
In the Bayesian framework, the aim is to infer the posterior distribution $p(\boldsymbol{\theta}\mid\mathbf{y})$ over the model parameters $\boldsymbol{\theta}$, given the observed data $\mathbf{y}$. From Bayes’ theorem, the posterior distribution can be expressed as:
$p(\boldsymbol{\theta}\mid\mathbf{y})=\frac{p(\mathbf{y}\mid\boldsymbol{\theta})p(\boldsymbol{\theta})}{p(\mathbf{y})},$ (1)
[…]
@@ 3.1 Conditional Prior VAE (CP-VAE)
The Conditional Prior Variational Autoencoder (CP-VAE) is an extension of the standard Conditional Variational Autoencoder (CVAE) (Sohn et al., 2015; Ivanov et al., 2019) specifically designed for simulation-based inference. We begin by introducing an auxiliary latent variable $\mathbf{z}$ to capture complex structures and dependencies within the joint distribution $p(\boldsymbol{\theta},\mathbf{y})$ of parameters $\boldsymbol{\theta}$ and simulated data $\mathbf{y}$. The latent variable $\mathbf{z}$ serves as a learned, low-dimensional summary statistic, efficiently encoding the most salient features of high-dimensional data relevant to characterize the posterior distribution. We formulate a latent variable model $p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ and approximate it using a variational distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$, parameterized by $\psi\in\Phi$, where $\Phi$ represents a family of distributions. The goal is to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true conditional distribution (Jordan et al., 1999; Kingma and Welling, 2014):
$\min_{\psi\in\Phi}D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\|p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\bigg{)}.$ (2)
@@
Using Bayes’ theorem and the chain rule of probability, the conditional distribution $p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ becomes:
$p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})=\frac{p(\boldsymbol{\theta}\mid\mathbf{z},\mathbf{y})p(\mathbf{z}\mid\mathbf{y})}{p(\boldsymbol{\theta}\mid\mathbf{y})}.$ (3)
@@
Substituting (3) into (2), and denoting the KL divergence $D_{KL}(q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\|p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta}))$ as $\mathcal{D}$, we obtain:
$\mathcal{D}$ $=\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})}\bigg{[}\log q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})-\log p(\mathbf{z}\mid\mathbf{y})$ (4)
$\qquad-\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})+\log p(\boldsymbol{\theta}\mid\mathbf{y})\bigg{]}.$
Rearranging the components of equation (4), we isolate $\log p(\boldsymbol{\theta}\mid\mathbf{y})$ to derive:
$\log p(\boldsymbol{\theta}\mid\mathbf{y})\geq$ $-D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\parallel p(\mathbf{z}\mid\mathbf{y})\bigg{)}$ (5)
$+\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})}\bigg{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})\bigg{]}.$
@@
This inequality stems from the fundamental non-negativity property of the KL divergence, establishing a lower bound on the log-posterior $p(\boldsymbol{\theta}\mid\mathbf{y})$. This bound forms the core of the optimization objective in the CP-VAE framework. Maximizing the log-posterior can be conveniently reformulated as minimizing its negative, which is a standard approach in optimization, leading to the following formulation:
$\mathcal{L}(\boldsymbol{\theta},\mathbf{y};\psi)=D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\,\|\,p(\mathbf{z}\mid\mathbf{y})\bigg{)}$ (6)
$-\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})}\bigg{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})\bigg{]}.$
@@
Figure 1: CP-VAE Architecture. The encoder $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ computes the parameters of the variational distribution for latent variables $\mathbf{z}$, conditioned on paired simulated data $\mathbf{y}$ and parameters $\boldsymbol{\theta}$. The prior network $p(\mathbf{z}\mid\mathbf{y})$ defines a data-dependent prior over $\mathbf{z}$, enforcing latent space regularization via the KL divergence. The decoder $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$ generates the distribution parameters for $\boldsymbol{\theta}$ based on $\mathbf{y}$ and $\mathbf{z}$. Both $\mathbf{z}$ and $\boldsymbol{\theta}$ are sampled using the reparameterization trick, ensuring end-to-end differentiability.
@@
The key advantage of this formulation lies in the prior distribution $p(\mathbf{z}\mid\mathbf{y})$ being conditioned on the observed data $\mathbf{y}$, allowing the model to adaptively shape its latent representation based on observed data, a crucial feature for effective amortized inference. This adaptive prior serves as a powerful regularization mechanism, ensuring that the latent space accurately captures the relevant features from the joint distribution $p(\boldsymbol{\theta},\mathbf{y})$. As a result, this approach delivers more precise posterior approximations, especially when dealing with data that exhibit complex conditional dependencies, which a fixed prior cannot adequately capture. Such adaptability is crucial for models aiming to generalize across a wide range of queries, making the CP-VAE framework well-suited for real-world applications with intricate data dependencies (Sohn et al., 2015).
The loss function presented in (6) is generally intractable due to the complexity of computing both the KL divergence between the variational distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ and the true conditional prior $p(\mathbf{z}\mid\mathbf{y})$, as well as the expectation over the latent variable $\mathbf{z}$. To make the optimization process tractable, we approximate the prior distribution $p(\mathbf{z}\mid\mathbf{y})$, the encoder distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$, and the decoder distribution $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$ using multivariate Gaussian distributions, each characterized by a mean vector and a covariance matrix. Since each dimension is assumed to be independent, we use a diagonal covariance matrix, which simplifies the calculation of KL divergence. The architecture of the model is illustrated in Figure 1. Detailed derivation of the loss function, along with further explanations of the underlying assumptions, are provided in Section 1 of the Appendix.
@@
To perform amortized inference with CP-VAE for a new observation $\mathbf{y}_{0}$, we first sample $N$ latent variables $\{\mathbf{z}^{(i)}\}_{i=1}^{N}$ from the conditional prior $p(\mathbf{z}\mid\mathbf{y}_{0})$ using the reparameterization trick (Kingma and Welling, 2014). For each latent variable $\mathbf{z}^{(i)}$, the decoder generates the mean and variance of the parameters, from which $\boldsymbol{\theta}^{(i)}$ is then sampled via reparameterization. By aggregating a large number of such samples, we can empirically approximate $p(\boldsymbol{\theta}\mid\mathbf{y}_{0})$, enabling statistical analysis and inference of $\boldsymbol{\theta}$ based on the observed data. The conditional prior, which is dependent on $\mathbf{y}_{0}$, allows the model to adapt its latent representation to the specific data, while the generative decoder maps latent variables to the parameter space, enabling CP-VAE to model complex, data-dependent relationships.
@@
Figure 2: UP-VAE architecture. The encoder $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ outputs the mean and variance of the approximate distribution, from which $\mathbf{z}$ is sampled (reparameterization trick). The KL divergence is computed between the encoder output and a standard normal prior. The theta decoder $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$ outputs the parameters for $\boldsymbol{\theta}$, and the data decoder $p(\mathbf{y}\mid\mathbf{z})$ outputs the parameters for $\mathbf{y}$ (reparameterization trick).

### v2

# Accelerating Simulation-Based Inference with Variational Autoencoders

@@ Abstract
We present a generative modeling approach based on the variational inference framework for likelihood-free simulation-based inference. The method leverages latent variables within variational autoencoders to efficiently estimate complex posterior distributions arising from stochastic simulations. We explore two variations of this approach distinguished by their treatment of the prior distribution. The first model adapts the prior based on observed data using a multivariate prior network, enhancing generalization across various posterior queries. In contrast, the second model utilizes a standard Gaussian prior, offering simplicity while still effectively capturing complex posterior distributions. We demonstrate the ability of the proposed approach to approximate complex posteriors while maintaining computational efficiency on well-established benchmark problems.
@@
The simulation-based inference (SBI) problem involves statistical inference of parameters $\bm{\theta}$ of a stochastic simulation model $M(\bm{\theta},\xi)$ containing random states $\xi$, from observed data $\bm{y}$ [Cranmer et al., 2020]. Once calibrated, the simulation model can be used to reason about, analyze and explain the observed data $\bm{y}$ in the context of the corresponding physical process . Several challenges arise in SBI owing to model stochasticity and potential multi-valuedness, where different sets of parameter values can produce similar observations, or similar parameters may lead to varied outputs. Consequently, the task is to infer the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$ in a Bayesian setting.
@@
A user-defined prior distribution $p(\bm{\theta})$ encodes domain knowledge or assumptions about plausible parameter values. The likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is however, often intractable due to intricate stochastic dynamics, high-dimensional integration, or latent variables making analytical computation impractical [Sisson et al., 2018]. Although a closed-form expression for the likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is unavailable, samples from this distribution can be generated by evaluating the simulator $M$ with varying $\xi$ for different $\bm{\theta}$ drawn from the prior, effectively yielding samples from the joint distribution $(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})$.
@@
Deep learning based approaches that directly approximate the model likelihood, or the posterior distribution have gained increasing popularity in the SBI setting [Zammit-Mangion et al., 2024]. Several approaches involve flow-based models, which are potentially compute-intensive and also restrict the design space of neural networks to invertible functions with computationally efficient Jacobian calculations [Rezende and Mohamed, 2015]. This constraint can limit the flexibility of the model in handling more complex data structures. Alternative approaches like Generative Adversarial Training for SBI (GATSBI) [Ramesh et al., 2022] use an adversarial network, where a generator produces simulations that mimic observed data, and a discriminator distinguishes between real and simulated data. However, the adversarial approach can suffer from training instability, mode collapse, and difficulties in balancing the generator and discriminator networks [Arjovsky and Bottou, 2022, Arjovsky et al., 2017], complicating training and limiting robustness. Recent approaches such as Simformer [Gloeckler et al., 2024] combine diffusion models with transformers and cross attention to enable highly accurate posterior estimation, but entail relatively high computational cost for training/inference.
@@
Contribution: We propose a variational inference approach for simulation-based inference, utilizing a Conditional Variational Autoencoder (C-VAE) architecture [Kingma and Welling, 2014, Sohn et al., 2015] that is simple, computationally efficient, interpretable and robust. As VAEs are not constrained by the invertibility requirement, they allow seamless integration of various neural network architectures without significant modifications as opposed to flow-based models. VAEs also avoid the instability associated with GANs due to adversarial learning dynamics. As a result, C-VAEs can handle complex data structures and dependencies more effectively, making them suitable for a wider range of tasks without requiring external summary networks [Cranmer et al., 2020]. The simplicity and efficiency also makes the approach particularly suitable for heavy workflows such as model exploration, which entail repeated solution of several inference problems.
@@
The paper is organized as follows. Section 1 outlines the related work. In Section 2, we provide a formal problem definition, followed by a detailed explanation of the proposed architecture in Sections 2.1 and 2.2. A comprehensive comparison on benchmark test problems is presented in Section 3. Section 4 discusses the results, while Section 5 concludes the paper by summarizing the key findings and further advancements.
@@ 1 Related Work
Recent deep learning advances have spurred a variety of SBI approaches. There exist methods that focus on approximating the likelihood directly [Hermans et al., 2020], while others concentrate on learning the posterior function [Papamakarios and Murray, 2016, Lueckmann et al., 2017, Greenberg et al., 2019]. A growing line of research pursues hybrid strategies that estimate both likelihood and posterior [Radev et al., 2023a, Glöckler et al., 2022, Wiqvist et al., 2021], aiming to balance flexibility and computational efficiency.
@@
Neural Posterior Estimation (NPE) [Papamakarios and Murray, 2016] was one of the earliest approaches towards one-step posterior estimation, introducing the use of conditional normalizing flows to approximate the posterior distribution. By incorporating techniques such as Neural Spline Flows (NSFs) and Mixture Density Networks (MDNs), NPE enhanced both the accuracy and scalability of simulation-based inference. In comparison to traditional approaches like Approximate Bayesian Computation (ABC) or Sequential Monte-Carlo (SMC)-ABC [Sisson et al., 2018] that rely on repeated simulations and rejection-based schemes, NPE and its derivatives such as Robust NPE [Ward et al., 2022] are more efficient and flexible.
@@
Sequential Neural Posterior Estimation (SNPE) improves NPE’s efficiency by employing sequential training, where initial parameters are sampled from the prior, and the model is iteratively refined by focusing on regions of the parameter space that are most relevant to the parameter inference task. This adaptive sampling allows for more efficient exploration of the posterior distribution. Early versions like SNPE-A [Papamakarios and Murray, 2016] and SNPE-B [Lueckmann et al., 2017] required correction steps to account for changes in the sampling distribution, adding complexity to the process. In contrast, Automatic Posterior Transformation (APT) or SNPE-C [Greenberg et al., 2019], eliminated this need, offering a more robust and streamlined approach by leveraging normalizing flows to refine the posterior iteratively. However, a limitation of sequential methods like APT is that they entail multiple rounds of training, extending the total training time.
@@
Jointly Amortized Neural Approximation (JANA) framework [Radev et al., 2023a] addresses posterior and likelihood estimation through joint amortization. JANA employs conditional invertible neural networks (cINNs) for both posterior and likelihood networks, allowing efficient transformations between the parameter space and latent variables. Additionally, JANA includes a trainable summary network sub-module optimized to extract maximally informative data representations in an end-to-end manner. Unlike previous approaches, JANA follows a fully amortized strategy, allowing the evaluation of normalized densities and generation of conditional random draws for parameter estimation and surrogate modeling.
The recently proposed Simformer [Gloeckler et al., 2024] employs a score-based diffusion model combined with transformer architectures and a cross-attention conditioning mechanism, allowing it to accurately learn complex, high-dimensional posteriors with substantial flexibility. However, diffusion-based approaches generally require iterative reverse-diffusion sampling, which, combined with the transformer’s computational overhead, results in slower amortized inference for higher-dimensional problems compared to other generative models.
@@
We compare our proposed approach with five popular SBI methods: JANA [Radev et al., 2023a], GATSBI [Ramesh et al., 2022], APT [Greenberg et al., 2019],NPE [Papamakarios and Murray, 2016] and Simformer [Gloeckler et al., 2024]. The motivation is to cover a variety of approaches (e.g., flow-based, adversarial, diffusion, one-shot, sequential) while noting that our focus is on efficiency and simplicity as opposed to solely the posterior approximation accuracy. We refer the reader to [Zammit-Mangion et al., 2024] and [Radev et al., 2023a] for further reading on deep learning methods for SBI.
@@ 2 Simulation Based Inference
In the Bayesian framework, the aim is to infer the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$ over the model parameters $\bm{\theta}$, given the observed data $\mathbf{y}$. From Bayes’ theorem, the posterior distribution can be expressed as:
$p(\bm{\theta}\mid\mathbf{y})=\frac{p(\mathbf{y}\mid\bm{\theta})p(\bm{\theta})}{p(\mathbf{y})},$ (1)
where $p(\mathbf{y})=\int p(\mathbf{y}\mid\bm{\theta})\,p(\bm{\theta})\,d\bm{\theta}$ is the marginal likelihood, which is a normalizing constant ensuring $p(\bm{\theta}\mid\mathbf{y})$ is a valid probability distribution. However, as $p(\mathbf{y})$ is independent of $\bm{\theta}$, it can be ignored in the computations involving the posterior.
@@ 2.1 Conditional Prior VAE (CP-VAE)
The Conditional Prior Variational Autoencoder (CP-VAE) extends the standard Conditional Variational Autoencoder (CVAE) [Sohn et al., 2015, Ivanov et al., 2019] specifically designed for simulation-based inference. We begin by introducing an auxiliary latent variable $\mathbf{z}$ to capture complex structures and dependencies within the joint distribution $p(\bm{\theta},\mathbf{y})$ of parameters $\bm{\theta}$ and simulated data $\mathbf{y}$. The latent variable $\mathbf{z}$ serves as a learned, low-dimensional summary statistic, efficiently encoding the most salient features of high-dimensional data relevant to characterize the posterior distribution. We formulate a latent variable model $p(\mathbf{z}\mid\mathbf{y},\bm{\theta})$ and approximate it using a variational distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})$, parameterized by $\psi\in\Phi$, where $\Phi$ represents a family of distributions. The goal is to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true conditional distribution [Jordan et al., 1999, Kingma and Welling, 2014]:
$\min_{\psi\in\Phi}D_{KL}\bigg(q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})\|p(\mathbf{z}\mid\mathbf{y},\bm{\theta})\bigg).$ (2)
@@
Using Bayes’ theorem and the chain rule of probability, the conditional distribution $p(\mathbf{z}\mid\mathbf{y},\bm{\theta})$ becomes:
$p(\mathbf{z}\mid\mathbf{y},\bm{\theta})=\frac{p(\bm{\theta}\mid\mathbf{z},\mathbf{y})p(\mathbf{z}\mid\mathbf{y})}{p(\bm{\theta}\mid\mathbf{y})}.$ (3)
@@
Substituting (3) into (2), and denoting the KL divergence $D_{KL}(q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})\|p(\mathbf{z}\mid\mathbf{y},\bm{\theta}))$ as $\mathcal{D}$, we obtain:
$\mathcal{D}$ $=\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})}\bigg[\log q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})-\log p(\mathbf{z}\mid\mathbf{y})$ (4)
$\qquad-\log p(\bm{\theta}\mid\mathbf{y},\mathbf{z})+\log p(\bm{\theta}\mid\mathbf{y})\bigg].$
Rearranging the components of equation (4), we isolate $\log p(\bm{\theta}\mid\mathbf{y})$ to derive:
$\log p(\bm{\theta}\mid\mathbf{y})\geq$ $-D_{KL}\bigg(q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})\parallel p(\mathbf{z}\mid\mathbf{y})\bigg)+\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})}\bigg[\log p(\bm{\theta}\mid\mathbf{y},\mathbf{z})\bigg].$ (5)
@@
This inequality stems from the fundamental non-negativity property of the KL divergence, establishing a lower bound on the log-posterior $p(\bm{\theta}\mid\mathbf{y})$. This bound forms the core of the optimization objective in the CP-VAE framework. Maximizing the log-posterior can be conveniently reformulated as minimizing its negative, which is a standard approach in optimization, leading to the following formulation:
$\mathcal{L}(\bm{\theta},\mathbf{y};\psi)=D_{KL}\bigg(q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})\,\|\,p(\mathbf{z}\mid\mathbf{y})\bigg)-\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})}\bigg[\log p(\bm{\theta}\mid\mathbf{y},\mathbf{z})\bigg].$ (6)
@@
Figure 1: CP-VAE Architecture. The encoder $q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})$ computes the parameters of the variational distribution for latent variables $\mathbf{z}$, conditioned on paired simulated data $\mathbf{y}$ and parameters $\bm{\theta}$. The prior network $p(\mathbf{z}\mid\mathbf{y})$ defines a data-dependent prior over $\mathbf{z}$, enforcing latent space regularization via the KL divergence. The decoder $p(\bm{\theta}\mid\mathbf{y},\mathbf{z})$ generates the distribution parameters for $\bm{\theta}$ based on $\mathbf{y}$ and $\mathbf{z}$. Both $\mathbf{z}$ and $\bm{\theta}$ are sampled using the reparameterization trick, ensuring end-to-end differentiability.
@@
The key advantage of this formulation lies in the prior distribution $p(\mathbf{z}\mid\mathbf{y})$ being conditioned on the observed data $\mathbf{y}$, allowing the model to adaptively shape its latent representation based on observed data, a crucial feature for effective amortized inference. This adaptive prior serves as a powerful regularization mechanism, ensuring that the latent space accurately captures the relevant features from the joint distribution $p(\bm{\theta},\mathbf{y})$. As a result, this approach delivers more precise posterior approximations, especially when dealing with data that exhibit complex conditional dependencies, which a fixed prior cannot adequately capture. Such adaptability is crucial for models aiming to generalize across a wide range of queries, making the CP-VAE framework well-suited for real-world applications with intricate data dependencies [Sohn et al., 2015].
The loss function in (6) is intractable due to the complexity of computing the KL divergence between the variational distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})$ and the conditional prior $p(\mathbf{z}\mid\mathbf{y})$, as well as the expectation over the latent variable $\mathbf{z}$. To address this, we approximate $p(\mathbf{z}\mid\mathbf{y})\sim\mathcal{N}(\bm{\mu}_{\mathbf{y}},\bm{\Sigma}_{\mathbf{y}})$, $q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})\sim\mathcal{N}(\bm{\mu}_{q},\bm{\Sigma}_{q})$, and $p(\bm{\theta}\mid\mathbf{y},\mathbf{z})\sim\mathcal{N}(\bm{\mu}_{\theta},\bm{\Sigma}_{\theta})$ as multivariate Gaussians with diagonal covariance matrices, where $\bm{\mu}_{\mathbf{y}}$, $\bm{\mu}_{q}$, and $\bm{\mu}_{\theta}$ are mean vectors, and $\bm{\Sigma}_{\mathbf{y}}$, $\bm{\Sigma}_{q}$, and $\bm{\Sigma}_{\theta}$ are covariance matrices. Here, $\mathbf{z}$ represents the latent variable with dimensionality $k_{\mathbf{z}}$. The final loss is:
$\mathcal{L}_{\text{CP-VAE}}$ $=\frac{1}{2}\bigg(\log\frac{|\bm{\Sigma}_{\mathbf{y}}|}{|\bm{\Sigma}_{q}|}-k_{\mathbf{z}}+\mathrm{tr}(\bm{\Sigma}_{\mathbf{y}}^{-1}\bm{\Sigma}_{q})+(\bm{\mu}_{\mathbf{y}}-\bm{\mu}_{q})^{\top}\bm{\Sigma}_{\mathbf{y}}^{-1}(\bm{\mu}_{\mathbf{y}}-\bm{\mu}_{q})\bigg)$ (7)
$\quad+\sum_{i=1}^{k}\log\sigma_{\theta_{i}}+\frac{1}{2N}\sum_{j=1}^{N}\sum_{i=1}^{k}\frac{(\theta_{i,j}-\mu_{\theta_{i}})^{2}}{\sigma_{\theta_{i}}^{2}},$
where $k$ is the dimensionality of $\bm{\theta}$. The model architecture is depicted in Figure 1.
@@
To perform amortized inference with CP-VAE for a new observation $\mathbf{y}_{0}$, we first sample $N$ latent variables $\{\mathbf{z}^{(i)}\}_{i=1}^{N}$ from the conditional prior $p(\mathbf{z}\mid\mathbf{y}_{0})$ using the reparameterization trick [Kingma and Welling, 2014]. For each latent variable $\mathbf{z}^{(i)}$, the decoder generates the mean and variance of the parameters, from which $\bm{\theta}^{(i)}$ is then sampled via reparameterization. By aggregating a large number of such samples, we can empirically approximate $p(\bm{\theta}\mid\mathbf{y}_{0})$, enabling statistical analysis and inference of $\bm{\theta}$ based on the observed data. The conditional prior, which is dependent on $\mathbf{y}_{0}$, allows the model to adapt its latent representation to the specific data, while the generative decoder maps latent variables to the parameter space, enabling CP-VAE to model complex, data-dependent relationships.
@@
Figure 2: UP-VAE architecture. The encoder $q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})$ outputs the mean and variance of the approximate distribution, from which $\mathbf{z}$ is sampled (reparameterization trick). The KL divergence is computed between the encoder output and a standard normal prior. The theta decoder $p(\bm{\theta}\mid\mathbf{y},\mathbf{z})$ outputs the parameters for $\bm{\theta}$, and the data decoder $p(\mathbf{y}\mid\mathbf{z})$ outputs the parameters for $\mathbf{y}$ (reparameterization trick).